OpenStudy (anonymous):

2b + 4 = -18 - 9b SHOW WORK PLEASE

OpenStudy (anonymous):

2b+4=-18-9b +9b +9b 11b+4=-18 -4 -4 11b=-22 /11 -11 B=2

OpenStudy (anonymous):

Go CheetahBooks1!